OpenStudy (anonymous):

Find the sine value of (7pi)/6

OpenStudy (idku):

7pi/6 = 210 degrees sin(210)=sin(180+30)=sin(180)cos(30)+cos(180)sin(30)= ... can you finish it off ?

OpenStudy (anonymous):

kinda... can you walk me through? cause i'm having a hard time reading what you wrote.. sorry I'm blind lol

OpenStudy (anonymous):

7pi/6 is in quadran 3 right? and ur reference angle is pi/6 . therefore the value is -1/2 since sine in q3 is negative

OpenStudy (anonymous):

get it?

OpenStudy (idku):

:)

OpenStudy (anonymous):

okay that makes more sense. but how do you get there?

OpenStudy (anonymous):

Like I understand how you got the reference angle but how did you get -1/2?

OpenStudy (nikato):

Do you have a unit circle with you?

OpenStudy (anonymous):

yes

OpenStudy (nikato):

Ok. Look at the coordinate for 7pi/6

OpenStudy (nikato):

You see that it's ( -sqrt3/2 , -1/2) And when were finding sin, it's asking for the y-coordinate which is -1/2

OpenStudy (nikato):

And if they ask for cos, it'll be the x-coordinate

OpenStudy (anonymous):

that makes so much more sense!! thanks so much!!

OpenStudy (nikato):

No problem. Glad you got it now

OpenStudy (anonymous):

glad u understand it :D took me a while to grasp trigonometry tooo. and always sine is the y axes and cosine is x axis just knowing that will come really handy in the future

OpenStudy (anonymous):

and now a plus to @nikato